can you please help me to solve this problem?
A typist would require 45 hours to type of manuscript. A second typist, who takes 2 min. more to each page, would require 60 hrs. How many pages of manuscript are there and how long per page does it take each typist?

Answer by Alan3354(69443)

A typist would require 45 hours to type of manuscript. A second typist, who takes 2 min. more to each page, would require 60 hrs. How many pages of manuscript are there and how long per page does it take each typist?
--------------------
The 1st typist takes 2700 minutes (45*60)
The 2nd takes 3600 minutes
-------------
n = # of pages
time per page t = n/2700 = (n+2)/3600
cross multiply
2700(n+2) = 3600n
2700n + 5400 = 3600n
900n = 5400
t = 6 min/page for the 1st typist
--> 8 min/page for the second
--------------------------------
n = 2700/6 = 450 pages